OpenStudy (anonymous):

assume a general exponential growth/decay model. A population of pigeons numbers 1,000 on the fifth day and 1,200 on the tenth day. What will it number on the twentieth day? a. 1,600 b. 1,728 c. 2,074 d. 1,546

OpenStudy (anonymous):

we can do this almost instantly \[P(t)=1000\times (1.2)^{\frac{t}{5}}\]

OpenStudy (anonymous):

the \(1.2\) from \(\frac{1200}{100}\) and the \(5\) i the denominator of the exponent because it grows by a factor of \(1.2\) in 5 days

OpenStudy (anonymous):

\[P(7)=1000\times (1.2)^{\frac{7}{5}}\] http://www.wolframalpha.com/input/?i=1000%281.2^%287%2F5%29%29